A pound of tin is worth $2 more than a pound of copper. Two pounds of tin are mixed with 8 pounds of copper to make bronze that sells for $4.85 per pound. How much is a pound of tin worth?
Answer by lwsshak3(11628)

A pound of tin is worth $2 more than a pound of copper. Two pounds of tin are mixed with 8 pounds of copper to make bronze that sells for $4.85 per pound. How much is a pound of tin worth?
..
let x=pound of tin worth($)
then x-2= pound of copper worth($)
2x+8(x-2)=10(4.85)
2x+8x-16=48.5
10x=48.5+16=64.5
x=$6.45 (copper)
x-2=$4.45 (bronze)
..
ans:
A pound of tin is worth=$6.45